find the displacement of plank 20 kg if a man of 60kg is walking on it
1
Expert's answer
2020-09-16T10:13:01-0400

solution

given data

mass of man (m)=60kg

mass of plank (M)=20 kg

assumption

length of plank is L meter and man starts to move one end of it to another end .

there are no friction between plank and surface.


so there are no any external force so center of mass of system will remain unchanged.

if plank move x meter distance then man will move (L-x) distance from one end to another.


com will not move because there are no any external force


"m(L-x)=Mx"


putting the value


"60(L-x)=20x"


"80x=60L"


"x=\frac{60L}{80}=0.75L"

so plank will move 0.75L meter.